Solve: dy/dx + y = y^3(cosx - sinx)
1
Expert's answer
2020-09-17T15:37:25-0400

Bernoulli differential equation


"\dfrac{dy}{dx}+y=(\cos x-\sin x)y^3, n=3"


Divide both sides by "y^3"


"\dfrac{1}{y^3}\dfrac{dy}{dx}+y^{-2}=\cos x-\sin x"

To find the solution, change the dependent variable from y to z, where "z=y^{1-3}=y^{-2}"


"\dfrac{dz}{dx}=(-\dfrac{2}{y^3})\dfrac{dy}{dx}"

"\dfrac{-2}{y^3}\dfrac{dy}{dx}-2y^{-2}=-2(\cos x-\sin x)"

"\dfrac{dz}{dx}-2z=-2\cos x+2\sin x"

"P(x)=-2"

We have the integrating factor


"I(x)=e^{\int(-2)dx}=e^{-2x}"

Then multiplying through by "I(x)", we get


"e^{-2x}\dfrac{dz}{dx}-2e^{-2x}z=2e^{-2x}(\sin x-\cos x)"

"\dfrac{d}{dx}(e^{-2x}z)=2e^{-2x}(\sin x-\cos x)"

"e^{-2x}z=2\int e^{-2x}(\sin x-\cos x)dx"

"\int e^{-2x}\sin xdx=-\cos x e^{-2x}-2\int e^{-2x}\cos xdx"

"\int e^{-2x}\cos xdx=\sin x e^{-2x}+2\int e^{-2x}\sin xdx"


"\int e^{-2x}\sin xdx=-\cos x e^{-2x}-2\sin x e^{-2x}-4\int e^{-2x}\sin xdx"

"5\int e^{-2x}\sin xdx=-\cos x e^{-2x}-2\sin x e^{-2x}"

"\int e^{-2x}\sin xdx=-\dfrac{1}{5}\cos x e^{-2x}-\dfrac{2}{5}\sin x e^{-2x}"

"\int e^{-2x}\cos xdx=\sin x e^{-2x}-\dfrac{2}{5}\cos x e^{-2x}-\dfrac{4}{5}\sin x e^{-2x}="

"=-\dfrac{2}{5}\cos x e^{-2x}+\dfrac{1}{5}\sin x e^{-2x}"


"e^{-2x}z=-\dfrac{2}{5}\cos x e^{-2x}-\dfrac{4}{5}\sin x e^{-2x}+"

"+\dfrac{4}{5}\cos x e^{-2x}-\dfrac{2}{5}\sin x e^{-2x}+C="

"=\dfrac{2}{5}\cos x e^{-2x}-\dfrac{6}{5}\sin x e^{-2x}+C"

"z=\dfrac{2}{5}\cos x -\dfrac{6}{5}\sin x +Ce^{2x}"

"\dfrac{1}{y^2}=\dfrac{2}{5}\cos x -\dfrac{6}{5}\sin x +Ce^{2x}"

"y^2=\dfrac{1}{\dfrac{2}{5}\cos x -\dfrac{6}{5}\sin x +Ce^{2x}}"